OpenStudy (anonymous):

The expression 2ln x -1/2ln(x^2-1) + ln(3-x^2) is the same as....

Nnesha (nnesha):

No. the expression isn't same as dots.

Nnesha (nnesha):

\[\large\rm \frac{ 2\ln (x-1) }{ 2 \ln(x^2-1) } +ln (3-x^2)\] like this ?

Nnesha (nnesha):

i'm not 100% sure but i think you have to apply ln rules quotient rule\[\large\rm ln y - ln x = \ln \frac{ x }{ y}\] subtraction to division product rule \[\large \rm ln x + \ln y = \ln ( x \times y )\] addition ----> multiplication power rule \[\large\rm ln x^y = y \ln x\]